# Sparse Training from Random Initialization: Aligning Lottery Ticket Masks using Weight Symmetry

### v1

@@ Abstract
The Lottery Ticket Hypothesis (LTH) suggests there exists a sparse LTH mask and weights that achieve the same generalization performance as the dense model while using significantly fewer parameters. However, finding a LTH solution is computationally expensive, and a LTH’s sparsity mask does not generalize to other random weight initializations. Recent work has suggested that neural networks trained from random initialization find solutions within the same basin modulo permutation, and proposes a method to align trained models within the same loss basin. We hypothesize that misalignment of basins is the reason why LTH masks do not generalize to new random initializations and propose permuting the LTH mask to align with the new optimization basin when performing sparse training from a different random init. We empirically show a significant increase in generalization when sparse training from random initialization with the permuted mask as compared to using the non-permuted LTH mask, on multiple datasets (CIFAR-10/100 & ImageNet) and models (VGG11 & ResNet20/50).
@@ 1 Introduction
Figure 1: Weight Symmetry and the Sparse Training Problem. A model with a single layer and only two parameters, $\mathbf{w}=({w_{0},w_{1}})$, operating on a single input scale $x_{0}$ has weight symmetry in the 2D loss landscape as illustrated above. In (1(a)) the original dense model, $\mathbf{w}_{A}$, is trained from a random dense initialization, $\mathbf{w}_{A}^{t=0}$ to a dense solution, $\mathbf{w}_{A}^{t=T}$, which is then pruned using weight magnitude resulting in the mask $\mathbf{m}_{A}=\left(1,0\right)$. In (1(b)), naively using the same mask to train a model, B, from a different random initialization will likely result in the initialization being far from a good solution. Permuting the mask to match the (symmetric) basin in which the new initialization is in will enable sparse training. See Figure 12 in Appendix E for the full figure also including the Lottery Ticket Hypothesis (LTH).
@@
    We hypothesize that LTH (Frankle & Carbin, 2019) fails to generalize well with a new random initialization due to a mismatch between the optimization basin of the winning ticket mask and the new random initialization’s solution basin. We propose a method based on permutation matching between two dense models, that permutes the winning ticket’s sparse mask to align with the optimization basin of the new random initialization. We empirically demonstrate on CIFAR-10/100 and ImageNet datasets using VGG11 and ResNet models of varying widths that permuting the LTH sparse mask to align with the new random initialization improves the performance of the trained model (permuted), compared to the model trained without permuting the sparse mask (naive).
@@ 4.1 Experimental Results
4.2 Diversity Analysis of Permuted Models
@@
Table 1: Ensemble Diversity Metrics for CIFAR-10/CIFAR-100. Although the mean test accuracy of LTH is higher, the ensemble of permuted models achieves better test accuracy due to better functional diversity of permuted models. Here we compare several measurements of function space similarity between the models including disagreement, which measures prediction differences  (Fort et al., 2020), and Kullback–Leibler (KL)/Jenson-Shannon (JS) divergence, which quantify how much the output distributions of different models differ  (Evci et al., 2022). As shown the permuted masks achieve similar diversity as computational expensive IMP solutions, also resulting in ensembles with a similar increase in generalization.
@@
Permutation matching is an NP-hard problem; the activation matching algorithm proposed by Ainsworth et al. (2023) does not find the global optimum; rather, it uses a greedy search to explore a restricted solution space. Therefore, in practice permutation matching does not perfectly align two models. However, it has been observed that for wider models, the algorithm can more closely align two models (Ainsworth et al., 2023; Sharma et al., 2024). To understand how the performance of the permuted model is affected by the approximation error of the matching algorithm, we evaluated the LMC and the accuracy of the permuted solution on ResNet20 models with varying layer widths. As shown in Figure 3, on increasing the layer width, the loss barrier of the interpolated network reduces, showing that permutation mapping becomes more accurate and aligns two models better. Also, it can be observed in Figures 5 and 6 that the permuted solution becomes close to the LTH solution on increasing the model width, showing that as the permutation matching becomes more accurate, the gap between the LTH and the permuted solution reduces.
@@
Figure 8: VGG11×$\times\{1\}$/CIFAR-10. Test accuracy of sparse network solutions at increasing rewind points for different sparsity levels. The dashed (- -) line shows the dense model accuracy. In  Figure 8(b), the permuted solution closely matches the LTH solution. However, beyond a certain rewind point, particularly for $k\geq 20$ the performance of the naive solution plateaus. Resulting in a more noticeable gap between the permuted and naive solutions compared to  Figure 8(a).
@@ A.1 Implementation Details for ResNet20 & VGG11 on CIFAR-10 and CIFAR-100
For convolutional neural networks, we train a modified version of the standard VGG11 implemented by (Jordan et al., 2023) on CIFAR-10. Primary differences are:
@@ A.2 Implementation Details for ResNet50 on ImageNet
We utilize the standard ResNet50 implementation provided by torchvision and customize PyTorch’s distributed data parallel codebase for training models on ImageNet (Contributors, 2024).
@@
For our set of experiments we used the ImageNet datatset (Deng et al., 2009). We apply the following standard data augmentation techniques to the training set:
@@ A.5 Additional Plots
Figure 10: Accuracy vs sparsity trend for ResNet20×$\times\{w\}$/CIFAR-100. Similar to the phenomenon seen in  Figure 9, with higher width the gap between permuted and naive solutions increases. As seen in  Figures 10(d), 10(h), 10(l) and 10(p) and the permuted solution approaches the LTH solution. The dashed (- -) line shows the dense model accuracy.
@@ Appendix B Early Matching
In our work, we train two dense models to convergence in order to find a permutation mapping using activation matching. However, it has been observed that the permutation mapping becomes stable early in the training (Sharma et al., 2024). In Table 15, we provide additional results on early matching with the CIFAR-10 dataset, which shows that models can be matched earlier in the training, thus reducing the computational cost of our method.
@@
In Section 4.2, IMP solution is trained independently over 5 different seeds with iterative pruning to obtain 5 different sparse/pruned solutions with different sparse masks/topologies ($M_{0}$, $M_{1}$, $M_{2}$, $M_{3}$, $M_{4}$, $M_{5}$). LTH ensemble is trained using the same mask ($M_{0}$) and init ($w_{0}$) over 5 different runs (with different data order). Random init $w_{0}$ defines the winning ticket for mask $M_{0}$. The permuted ensemble is trained using 5 different permutations ($\pi_{1}$, $\pi_{2}$, $\pi_{3}$, $\pi_{4}$, $\pi_{5}$) of the same mask ($M_{0}$) with five different random inits ($w_{1}$, $w_{2}$, $w_{3}$, $w_{4}$, $w_{5}$).

### v2

@@ Abstract
The Lottery Ticket Hypothesis (LTH) suggests there exists a sparse LTH mask and weights that achieve the same generalization performance as the dense model while using significantly fewer parameters. However, finding a LTH solution is computationally expensive, and a LTH’s sparsity mask does not generalize to other random weight initializations. Recent work has suggested that neural networks trained from random initialization find solutions within the same basin modulo permutation, and proposes a method to align trained models within the same loss basin. We hypothesize that misalignment of basins is the reason why LTH masks do not generalize to new random initializations and propose permuting the LTH mask to align with the new optimization basin when performing sparse training from a different random init. We empirically show a significant increase in generalization when sparse training from random initialization with the permuted mask as compared to using the non-permuted LTH mask, on multiple datasets (CIFAR-10/100 & ImageNet) and models (VGG11 & ResNet20/50). Our codebase for reproducing the results is publicly available at here.
@@ 1 Introduction
Figure 1: Weight Symmetry and the Sparse Training Problem. A model with a single layer and only two parameters, $\mathbf{w}=({w_{0},w_{1}})$, operating on a single input $x_{0}$ has weight symmetry in the 2D loss landscape as illustrated above. In (1(a)) the original dense model, $\mathbf{w}_{A}$, is trained from a random dense initialization, $\mathbf{w}_{A}^{t=0}$ to a dense solution, $\mathbf{w}_{A}^{t=T}$, which is then pruned using weight magnitude resulting in the mask $\mathbf{m}_{A}=\left(1,0\right)$. In (1(b)), naively using the same mask to train a model, B, from a different random initialization will likely result in the initialization being far from a good solution. Permuting the mask to match the (symmetric) basin in which the new initialization is in will enable sparse training. See Figure 12 in Appendix E for the full figure also including the Lottery Ticket Hypothesis (LTH).
@@
    We hypothesize that the LTH (Frankle & Carbin, 2019) fails to generalize well with a new random initialization due to a mismatch between the optimization basin of the winning ticket mask and the new random initialization’s solution basin. We propose a method based on permutation matching between two dense models, that permutes the winning ticket’s sparse mask to align with the optimization basin of the new random initialization. We empirically demonstrate on CIFAR-10/100 and ImageNet datasets using VGG11 and ResNet models of varying widths that permuting the LTH sparse mask to align with the new random initialization improves the performance of the trained model (permuted), compared to the model trained without permuting the sparse mask (naive).
@@ 4 Results
4.2 Diversity Analysis of Permuted Models.
@@
Table 1: Ensemble Diversity Metrics for CIFAR-10/CIFAR-100. Although the mean test accuracy of LTH is higher, the ensemble of permuted models achieves better test accuracy due to better functional diversity of permuted models. Here we compare several measurements of function space similarity between the models including disagreement, which measures prediction differences  (Fort et al., 2020), and Kullback–Leibler (KL)/Jenson-Shannon (JS) divergence, which quantify how much the output distributions of different models differ  (Evci et al., 2022). As shown, the permuted masks achieve similar diversity as computational expensive IMP solutions, also resulting in ensembles with a similar increase in generalization.
@@
Permutation matching is an NP-hard problem; the activation matching algorithm proposed by Ainsworth et al. (2023) does not find the global optimum; rather, it uses a greedy search to explore a restricted solution space. Therefore, in practice, permutation matching does not perfectly align two models. However, it has been observed that for wider models, the algorithm can more closely align two models (Ainsworth et al., 2023; Sharma et al., 2024). To understand how the performance of the permuted model is affected by the approximation error of the matching algorithm, we evaluated the LMC and the accuracy of the permuted solution on ResNet20 models with varying layer widths. As shown in Figure 3, on increasing the layer width, the loss barrier of the interpolated network reduces, showing that permutation mapping becomes more accurate and aligns two models better. Also, it can be observed in Figures 5 and 6 that the permuted solution becomes close to the LTH solution on increasing the model width, showing that as the permutation matching becomes more accurate, the gap between the LTH and the permuted solution reduces.
@@
Figure 8: VGG11×$\times\{1\}$/CIFAR-10. Test accuracy of sparse solutions at increasing rewind points for different sparsity levels. The dashed (- -) line shows the dense model accuracy. In  Figure 8(b), the permuted solution closely matches the LTH solution. However, beyond a certain rewind point, i.e. for $k\geq 20$ the performance of the naive solution plateaus. Resulting in a more noticeable gap between the permuted and naive solutions compared to  Figure 8(a).
@@ Acknowledgements
Contribution Statement
All authors contributed to the writing of the paper. MA and RJ implemented the method, conducted most experiments, and contributed the majority of the writing. MA and YI played key roles in designing the methodology and experimental setup. ES contributed to the design of the experimental setup, provided key insights, hypothesized the connection to the conclusion drawn by Paul et al. (2023), as summarized in Figure 4 and our contributions, and ran experiments to validate it. RGK and YI, as senior authors, provided feedback on the writing and methodology. YI conceptualized the research idea, designed Figures 1, 2 and 12, and also contributed feedback on code and experiments throughout the project.
@@ A.1 Implementation Details for ResNet20 & VGG11 on CIFAR-10 and CIFAR-100
For convolutional neural networks, we train a modified version of the standard VGG11 implemented by Jordan et al. (2023) on CIFAR-10. Primary differences are:
@@ A.2 Implementation Details for ResNet50 on ImageNet
We utilize the standard ResNet50 implementation provided by torchvision and customize PyTorch’s distributed data parallel codebase for training models on ImageNet (Paszke et al., 2019).
@@
For our set of experiments we used the ImageNet dataset (Deng et al., 2009). We apply the following standard data augmentation techniques to the training set:
@@ A.5 Additional Plots
Figure 10: Accuracy vs sparsity trend for ResNet20×$\times\{w\}$/CIFAR-100. Similar to the phenomenon seen in  Figure 9, with higher width, the gap between permuted and naive solutions increases. As seen in  Figures 10(d), 10(h), 10(l) and 10(p) and the permuted solution approaches the LTH solution. The dashed (- -) line shows the dense model accuracy.
@@ Appendix B Early Matching
In our current methodology, we train both models $A$ and $B$ to convergence, resulting in the weight configurations $\textbf{w}_{A}^{t=T}$ and $\textbf{w}_{B}^{t=T}$, respectively. However, it has been observed in (Sharma et al., 2024) that it is possible to find permutation mapping earlier in training. In effort to reduce the computational cost associated with our approach, we aim to find a permutation $\pi$ that allows us to suitably align the weights of model $A$ at convergence, $\textbf{w}_{A}^{t=T}$, with the weights of model $B$ at an earlier training iteration $i\ll T$, $\textbf{w}_{B}^{t=i}$. In Table 15, we provide additional results on early matching with the CIFAR-10 dataset, which shows that models can be matched earlier in the training, thus reducing the computational cost of our method.
@@
In Section 4.2, the IMP solution is trained independently over $5$ different seeds with iterative pruning to obtain $5$ different sparse/pruned solutions with different sparse masks/topologies ($M_{1}$, $M_{2}$, $M_{3}$, $M_{4}$, $M_{5}$). The LTH ensemble is trained using the same mask ($M_{1}$) and initialization ($w_{1}$) over $5$ different runs (with different data order). Random initialization $w_{1}$ defines the winning ticket for mask $M_{1}$. The permuted ensemble is trained using $5$ different permutations ($\pi_{1}$, $\pi_{2}$, $\pi_{3}$, $\pi_{4}$, $\pi_{5}$) of the same mask ($M_{1}$) with five different random weight initializations ($w_{1}$, $w_{2}$, $w_{3}$, $w_{4}$, $w_{5}$).